If the difference between the compound interest,


compounded every six months, and:the simple interest


on a certain sum of money at the rate of 12% per annum


for one year is Rs. 36, the sum is:


(A) Rs.lO,OOO (B) Rs . 12,000 (C) Rs. 15,000 (D) Rs. 18,000

Grade:Upto college level

2 Answers

Pratik Tibrewal
10 years ago
Let the money be Rs x
If it is simple interest, then the Interest is: x * 12/100
if compound Interest.: then interest = x(1+6/100)^2 - x = x ( 1.06*1.06 - 1)
Difference in interest :
x(1.06 *1.06 - 1 - 12/100) = 36
x(0.0036) = 36
=> x = 10,000
(A)

Thanks and Regards,
Pratik Tibrewal,
askiitians faculty,
BTech
Pratik Tibrewal
10 years ago
S.I = x*12/100
C.I = x( 1.06*1.06 - 1)
C.I. - S.I. = x(1.06 *1.06 - 1 -0.12) = x(0.0036) = 36
=> x = 10000
A)

Thanks and Regards,
Pratik Tibrewal,
askiitians faculty,
BTech